A 5% acid solution is mixed with a 9% acid solution. How many ounces of the 5% acid solution must you mix with the 10 ounces of 9% acid solution to obtain 40ounces of 6%acid solution?
Answer by stanbon(75887)

A 5% acid solution is mixed with a 9% acid solution. How many ounces of the 5% acid solution must you mix with the 10 ounces of 9% acid solution to obtain 40ounces of 6%acid solution?
----------
Acid EQUATION: 0.05x + 0.09*10 = 0.06*40
0.05x + 0.9 = 2.4
0.05x = 1.5
5x = 150
x = 30 oz(amount of 5% that must be added)
================
Cheers,
Stan H.